# London WC1E 6BT, United Kingdom

### v1

@@ 4.2 Cosmic Microwave Background
After the end of BBN, any particle that injects energy in the primordial plasma will modify, to some extent, the observations of the CMB. The impact of energy injected before recombination would work to heat up the plasma and hence cause changes to the observed anisotropies in the CMB power spectrum. However, since the major source of energy injected is via ALP decays to neutrinos, the energy injected into the photon bath is negligible, and thus we do not expect constraints arising from CMB anisotropies. Late time decays of ALPs could also cause spectral distortions in the black body spectrum of the CMB if the decays heat up the photon bath significantly Balazs:2022tjl . Fortunately, as before, the decay rate from ALPs to photons is negligible and we assume that secondary energy injected into the photon spectrum from the decays of ALPs can be safely ignored.
The most significant impact of the decays of ALPs for recombination is the modification of the neutrino temperature. Since neutrino decoupling, the photon and neutrino temperatures evolved independently which, even in $\Lambda$CDM, causes a value of the effective neutrino degrees of freedom as $N_{\rm eff}=3.044\pm 0.384$ Planck:2018vyg . ALP decays before the formation of the CMB increase the neutrino temperature, as in eq.(38). This increase of the temperature of the neutrino bath with respect to that of the photon bath, causes the value of $N_{\rm eff}$ to increase from the expectations of $\Lambda$CDM, as
$N_{\rm eff}^{\rm CMB}=N_{\rm eff}^{\rm BBN}\left(\frac{11}{4}\right)^{\tfrac{4}{3}}\left(\frac{T_{\nu}}{T_{\gamma}}\right)^{4}$ (40)
@@ 4.3 SN1897A
The core of supernovae (SN) are very hot and dense systems, with temperatures of the order of $T\sim 30$ MeV. Most of the particles created in such energetic medium are trapped and contribute to the energy transfer inside the SN core. Weakly coupled particles, however, can free-stream and escape the core, contributing to the cooling of the SN. The primary source of cooling for SNs are neutrinos, which can escape as long as they have energies $E_{\nu}\lesssim 30$ MeV. This neutrino burst was observed for SN1987A by various water Cherenkov detectors, including Kamiokande-II Kamiokande-II:1987idp ; Totsuka:1988iyh . Besides neutrinos, the HNLs and ALPs in our model could also escape the SN core as they interact very weakly with the SN inner medium, and contribute to SN cooling Dolgov:2000pj ; Fuller:2008erj ; Mastrototaro:2019vug ; Fiorillo:2022cdq ; Syvolap:2023trc ; Akita:2023iwq . This additional source of cooling is very constrained by measurements of the luminosity of SN187A. Secondary decays of HNLs and ALPs into neutrinos produce a high-energetic additional flux of active neutrinos that could have been detected alongside the normal neutrino burst Fiorillo:2022cdq ; Syvolap:2023trc . The combination of SN cooling and the additional neutrino flux can impose strong constraints on our model for high couplings.
@@
For both HNLs and ALPs, the constraint on the secondary neutrino flux is much stronger than the constraint from SN cooling Syvolap:2023trc ; Akita:2023iwq , and hence we only include the former in our study. Typical constraints on the secondary neutrino flux from HNL decays assume decay rates for HNLs in the absence of additional channels, so the $N\to\pi\nu$ and $N\to 3\nu$ decays dominates Mastrototaro:2019vug . In our model the primary decay channel for HNLs is to ALPs, which consecutively decay to a pair of neutrinos. Hence, neglecting resonant effects, the secondary neutrino flux from HNLs can be approximated to that arising from the 3-body decay of the HNLs to neutrinos. Even though the 2-body decay branching ratio is around 7 times larger than the 3-body decays for HNL masses in the 100 MeV range, Cherenkov detectors are more sensitive to the 3-body decays, since the interaction cross section of antineutrinos, only produced in 3-body decays, is around 100 times larger than that of neutrinos Kamiokande-II:1987idp . Therefore, SN limits on traditional HNL decays can be directly applied to our scenario. We thus use the limits on HNL masses and mixing from Fig. 2 of Akita:2023iwq to constrain the large mixing angle regions in our model.
@@
In addition to HNLs, ALPs can also be produced in the core of SN. The expected secondary neutrino flux from decaying ALPs with masses below the keV scale is smaller than that from HNL decays that can escape the SN core. However, heavy HNLs cannot escape and thus SN constraints from the production of ALPs are stronger for $m_{N}\gtrsim 400$ MeV. From Fig. 5 of Akita:2023iwq , we can see that the coupling between the ALP and the electron-neutrino $g_{ae}$ must be $g_{ae}\lesssim 10^{-7}$ for keV-scale ALPs. In our model, $g_{ae}$ is given by Eq. (26), which imposes only weak limits for heavy HNLs and large HNL-ALP couplings, beyond the ranges considered in our study.
@@
Lastly, there could be additional constraints from secondary decays of HNLs and/or ALPs to photons Hoof:2022xbe . However, these are negligible as the photonic branching ratios of both particles are very small for the masses considered.
@@ 4.4 Other astrophysical constraints
In addition to the astrophysical constraints from SN cooling and its secondary fluxes, there is a plethora of possible astrophysical probes of ALPs. The production of ALP in the core of white dwarfs or RGB stars can lead to strong cooling, and even provide an explanation for the observed cooling hints Giannotti:2015kwo ; DiLuzio:2020wdo . In fact, it has been shown that an ALP with a coupling to electrons of the order $g_{ae}\sim 10^{-13}$ provides a good fit to stellar cooling data DiLuzio:2020wdo . However, in our model the ALP-electron coupling is very small, so much that the value $g_{ae}\sim 10^{-13}$ can only be reached for mixing $|U_{eN}|^{2}\sim 10^{-1}$ and HNL masses $m_{N}\sim 1$ TeV. Since this parameter region is beyond our scope, and anyways disfavoured by various cosmological constraints, as we will see later, we neglect the constraints from stellar cooling in our study.
Long-lived ALPs that survive after recombination may still be observable today through their decay products. The photons injected by ALP decays are the most detectable candidates, as they can be probed in observations of the extra-galactic background light (EBL) or via X-rays Cadamuro:2011fd . For ALP masses below the keV scale, the constraints on the ALP-photon coupling from X-ray and EBL surveys require that $g_{a\gamma}\lesssim 10^{-15}$ GeV-1 Cadamuro:2011fd ; Balazs:2022tjl . In our model, however, the ALP-photon coupling is derived at two-loop order, and therefore extremely small. As above, only for very large mixing $U_{eN}^{2}\sim 10^{-3}$ and large HNL masses $m_{N}\sim 1$ TeV, has this constraint any effect. Consequently, we also ignore any astrophysical constraints from late-time photonic decays of ALPs.
@@ 4.5 Impact on direct HNL searches
Future HNL searches, especially those based on long-lived signatures, will probe small active-sterile mixing strengths approaching the seesaw expectation. Prominent examples of proposed and planned searches are PIONEER PIONEER:2022yag , NA62 Dias:2022fwf ; Drewes:2018gkc , DUNE Ballett:2019bgd , SHiP SHiP:2018xqw , FCC-ee Blondel:2022qqo . A dedicated analysis is needed to present a study of the sensitivity in our scenario depending on the specific search strategy. Such a detailed analysis is beyond the scope of this paper and will be considered as future work. In this section, we will assess how the sensitivity is modified in presence of new HNL-ALP coupling in the context of DUNE. We qualitatively describe the approach to calculate the expected number of signal events in the DUNE near detector (ND) based on the analysis Bolton:2022tds . The ND is located at a distance of $L=574$ m from the HNL production point, with a transverse cross-section of $A=12~{}\rm{m}^{2}$ and a depth of $\Delta L=5$ m along the beam axis. Following Bolton:2022tds , we write the expected number of signal events as
$N_{\rm{sig}}=N_{P}\times\rm{Br}(P\rightarrow N)\times\rm{Br}(N\rightarrow\rm{charged})\times\epsilon_{\rm{geo}},$ (41)
@@
$\epsilon_{\rm{geo}}=e^{-\frac{m_{N}\Gamma_{N}}{p_{N_{z}}}L}\left(1-e^{-\frac{m_{N}\Gamma_{N}}{p_{N_{z}}}\Delta L}\right).$ (42)
Here, $p_{N_{z}}$ is the momentum of the HNL along the beam axis in the lab frame, where we have considered $p_{N_{z}}=7.5$ GeV, following the simulation of meson production from a $pp$ collision at $\sqrt{s}=15$ GeV Bolton:2022tds . As we have not explicitly simulated any events and analysed with respect to detector level cuts, for simplicity we will consider, in our analysis, that all produced HNL events will be accepted at detector level.
@@
$\frac{N^{\prime}_{\rm{sig}}}{N_{\rm{sig}}}=\frac{\rm{Br^{\prime}}(N\rightarrow\rm{charged})}{\rm{Br}(N\rightarrow\rm{charged})}\frac{\epsilon^{\prime}_{\rm{geo}}}{\epsilon_{\rm{geo}}}$ (43)
@@
$\frac{\epsilon^{\prime}_{\rm{geo}}}{\epsilon_{\rm{geo}}}=\exp\left[-\frac{m_{N}}{p_{N_{z}}}\Gamma(N\to a\nu)L\right]\frac{\Gamma^{\prime}_{N}}{\Gamma_{N}},$ (44)
for a shallow detector depth $\Delta L$, i.e., for $(m_{N}/p_{N_{z}})\Gamma^{\prime}_{N}\Delta L\ll 1$. The ratio of decay rates in this expression is equal to $\Gamma^{\prime}_{N}/\Gamma_{N}=\text{Br}(N\to\text{charged})/\rm{Br^{\prime}}(N\to\text{charged})$, thus cancelling the corresponding ratio in Eq. (43) in this limit. We thus have
$\frac{N^{\prime}_{\rm{sig}}}{N_{\rm{sig}}}=\exp\left[-\frac{m_{N}}{p_{N_{z}}}\Gamma(N\to a\nu)L\right],$ (45)
@@
As another example for a direct search experiment we consider the existing NA62 experiment which uses a different search strategy as compared to DUNE. The NA62 experiment Dias:2022fwf used a secondary 75 GeV hadron beam containing a fraction of kaons, and has been able to probe the decays $K^{+}\to\ell^{+}N$. For small active-sterile mixing the decay length of HNL is much larger than the 75 m detector size and the process is characterised by a single detected track, that of the charged lepton – a positive signal is a peak in its missing mass distribution. As this experiment is currently insensitive to the decay of the HNLs, the presence of the $N\to a\nu$ decay channel will not affect the search as long as the decay length remains large.
@@ 5 Results and discussion
The aim of this work is to find valid scenarios where the traditional BBN bound on HNLs Boyarsky:2020dzc is relaxed due to the primary decays to ALPs. For this purpose we have performed some small scale parameter scans around the benchmark scenarios in Table 1 to illustrate this effect [^3] We performed a grid scan on two parameters of the model, $m_{N}$ and $|U_{eN}|^{2}$. This simplification is enough for our purposes as we only intend to highlight the crucial features of the model, but do not attempt to map the available parameter space nor perform any statistical interpretation of the results. For more details on the rigorous treatment of inference and statistics in physics see AbdusSalam:2020rdj .. Figure 5 shows the results of these parameter scans, varying the values of $m_{N}=[10^{-3},10^{2}]$ GeV and $|U_{eN}|^{2}=[10^{-13},10^{-6}]$. The top row corresponds to scans around scenarios 1 and 2, with only $m_{a}$ and $f_{a}$ fixed, and, where the blue diamond markers are the fixed values of $m_{N}$ and $|U_{eN}|^{2}$ from Table 1. The bottom row shows scans of scenarios 3 and 4, which have a lighter ALP mass of $m_{a}=10$ eV.
@@
The strongest cosmological constraint, shown as the shaded brown region in the panels, arises from the increase on the neutrino temperature due to ALP decays, which shows as a modification of the $N_{\rm eff}$ at the time of recombination. $N_{\rm eff}$ constrains the larger HNL masses and mixing, corresponding to shorter ALP lifetimes. Consequently, ALPs decaying before the onset of BBN, or shortly after, would cause a sufficient increase of the neutrino temperature and be excluded by the Planck measurement of the $N_{\rm eff}$ value. Smaller HNL-ALP couplings cause longer ALP lifetimes, and thus the right-handed panels with larger $f_{a}$ have weaker $N_{\rm eff}$ constraints. However, for the bottom panels with lower $m_{a}$, even though the lifetime is also longer and one would expect the limit to be weaken, the opposite effect occurs and the $N_{\rm eff}$ limit is significantly stronger. This happens because lighter ALPs are ultra-relativistic for a longer period between BBN and recombination, and thus have a stronger effect on the expansion rate, and thus on the neutrino temperature and value of $N_{\rm eff}$. Other cosmological constraints such as, energy injection before BBN or the relic abundance of ALPs are subleading and thus not shown. The region above the purple contour is, in principle, excluded due to the astrophysical constraints from supernova SN1987A, arising from the modified neutrino flux due to the production of HNLs in the core of the supernova. Nevertheless, the uncertainties in the calculation of the SN neutrino flux are high and depend on the choice of SN core model Carenza:2023old , hence we show in purple a conservative expectation of the limit, but refrain from excluding models only on the basis of this SN constraint. Supernova constraints due to ALP production, which is largely independent of $m_{a}$ in the ranges of interest, only apply for large values of $|U_{eN}|^{2}$ and $m_{N}$, which however resides well outside of our considered region.
@@
In Fig. 6 we present this newly available parameter region which may be the target of future dedicated searches for HNLs. In the left panel, we show the corresponding parameter space for scenario 1 ($f_{a}=1$ TeV, $m_{a}=1$ keV) as described in previous section, while the right panel corresponds to the scenario 2 ($f_{a}=10^{2.5}$ TeV, $m_{a}=1$ keV). The gray shaded regions in both panels are disfavored from various cosmological and astrophysical constraints discussed in Sec. 4, and shown coloured in Figure 5. The gray dashed curve denotes the BBN bound from the HNL only decaying to SM decay channels i.e., in absence of the ALP. The corresponding seesaw regime is shown in blue considering $1~{}\rm{meV}~{}\lesssim m_{\nu}\lesssim 100~{}\rm{meV}$. The red dashed contour corresponds to the future sensitivity of DUNE HNL searches in standard HNL models (without ALPs), while the solid red contours in both panels denote the resultant sensitivity contours for observation of 6 events in presence of the new $N\rightarrow a\nu$ decay channel in respective benchmark scenarios with two different $f_{a}$ values. Additionally, as described in Tab. 1, in blue diamonds two benchmark scenarios with fixed HNL mass and mixing $m_{N}=10^{-1}$ GeV, $|U_{eN}|^{2}=10^{-10}$ and $m_{N}=10^{-0.4}$ GeV, $|U_{eN}|^{2}=10^{-9.2}$, have been shown. The brown dashed contour denotes the present exclusion contour from NA62 experiment. Furthermore, considering the discussion of subsection 4.5, we have also highlighted in brown the region where the approximation holds (with $L_{\rm{NA}62}$ = 75 m and $p_{N_{z}}\sim$ 30 GeV as the $K^{+}$ beams have energy $\sim 75$ GeV Dias:2022fwf ), the present sensitivity contour remains unaffected by the new axion decay channel of HNL in this region. From Figure 6 it is evident that the available regions constrained by several astrophysical and cosmological constraints can be well probed in the future by DUNE and NA62 in presence of this new decay channel of HNLs.

### v2

@@ 4.2 Cosmic Microwave Background
After the end of BBN, any particle that injects energy in the primordial plasma will modify, to some extent, the observations of the CMB. The impact of energy injected before recombination would work to heat up the plasma and hence cause changes to the observed anisotropies in the CMB power spectrum. However, since the major source of energy injected is via ALP decays to neutrinos, the energy injected into the photon bath is negligible, and thus we do not expect constraints arising from CMB anisotropies. Late time decays of ALPs could also cause spectral distortions in the black body spectrum of the CMB if the decays heat up the photon bath significantly [44]. Fortunately, as before, the decay rate from ALPs to photons is negligible and we assume that secondary energy injected into the photon spectrum from the decays of ALPs can be safely ignored.
The most significant impact of the decays of ALPs for recombination is the modification of the neutrino temperature. Since neutrino decoupling, the photon and neutrino temperatures evolved independently which, even in $\Lambda$CDM, causes a value of the effective neutrino degrees of freedom as $N_{\rm eff}=3.044\pm 0.384$ [61]. ALP decays before the formation of the CMB increase the neutrino temperature, as in eq.(3.11). This increase of the temperature of the neutrino bath with respect to that of the photon bath, causes the value of $N_{\rm eff}$ to increase from the expectations of $\Lambda$CDM, as
$N_{\rm eff}^{\rm CMB}=N_{\rm eff}^{\rm BBN}\left(\frac{11}{4}\right)^{\tfrac{4}{3}}\left(\frac{T_{\nu}}{T_{\gamma}}\right)^{4}$ (4.2)
@@ 4.3 SN1987A
The core of supernovae (SN) are very hot and dense systems, with temperatures of the order of $T\sim 30$ MeV. Most of the particles created in such energetic medium are trapped and contribute to the energy transfer inside the SN core. Weakly coupled particles, however, can free-stream and escape the core, contributing to the cooling of the SN. The primary source of cooling for SNs are neutrinos, which can escape as long as they have energies $E_{\nu}\lesssim 30$ MeV. This neutrino burst was observed for SN1987A by various water Cherenkov detectors, including Kamiokande-II [63, 64]. Besides neutrinos, the HNLs and ALPs in our model could also escape the SN core as they interact very weakly with the SN inner medium, and contribute to SN cooling [65, 66, 67, 68, 69, 70]. This additional source of cooling is very constrained by measurements of the luminosity of SN1987A. Secondary decays of HNLs and ALPs into neutrinos produce a high-energetic additional flux of active neutrinos that could have been detected alongside the normal neutrino burst [68, 69]. The combination of SN cooling and the additional neutrino flux can impose strong constraints on our model for high couplings.
@@
For both HNLs and ALPs, the constraint on the secondary neutrino flux is much stronger than the constraint from SN cooling [69, 70], and hence we only include the former in our study. Typical constraints on the secondary neutrino flux from HNL decays assume decay rates for HNLs in the absence of additional channels, so the $N\to\pi\nu$ and $N\to 3\nu$ decays dominates [67]. Other studies involving SN constraints on lighter HNLs can be found in [71, 72, 65, 73]. In our model the primary decay channel for HNLs is to ALPs, which consecutively decay to a pair of neutrinos. Hence, neglecting resonant effects, the secondary neutrino flux from HNLs can be approximated to that arising from the 3-body decay of the HNLs to neutrinos. Even though the 2-body decay branching ratio is around 7 times larger than the 3-body decays for HNL masses in the 100 MeV range, Cherenkov detectors are more sensitive to the 3-body decays, since the interaction cross section of antineutrinos, only produced in 3-body decays, is around 100 times larger than that of neutrinos [63]. Therefore, SN limits on traditional HNL decays can be directly applied to our scenario. We thus use the limits on HNL masses and mixing from Fig. 2 of [70] to constrain the large mixing angle regions in our model.
@@
In addition to HNLs, ALPs can also be produced in the core of SN. The expected secondary neutrino flux from decaying ALPs with masses below the keV scale is smaller than that from HNL decays that can escape the SN core. However, heavy HNLs cannot escape and thus SN constraints from the production of ALPs are stronger for $m_{N}\gtrsim 400$ MeV. From Fig. 5 of [70], we can see that the coupling between the ALP and the electron-neutrino $g_{ae}$ must be $g_{ae}\lesssim 10^{-7}$ for keV-scale ALPs. In our model, $g_{ae}$ is given by Eq. (2.26), which imposes only weak limits for heavy HNLs and large HNL-ALP couplings, beyond the ranges considered in our study.
@@
Lastly, there could be additional constraints from secondary decays of HNLs and/or ALPs to photons [74, 75, 76, 77, 78, 79]. However, these are negligible as the photonic branching ratios of both particles are very small for the masses considered.
@@ 4.4 Other astrophysical constraints
In addition to the astrophysical constraints from SN cooling and its secondary fluxes, there is a plethora of possible astrophysical probes of ALPs. The production of ALP in the core of white dwarfs or RGB stars can lead to strong cooling, and even provide an explanation for the observed cooling hints [80, 81]. In fact, it has been shown that an ALP with a coupling to electrons of the order $g_{ae}\sim 10^{-13}$ provides a good fit to stellar cooling data [81]. However, in our model the ALP-electron coupling is very small, so much that the value $g_{ae}\sim 10^{-13}$ can only be reached for mixing $|U_{eN}|^{2}\sim 10^{-1}$ and HNL masses $m_{N}\sim 1$ TeV. Since this parameter region is beyond our scope, and anyways disfavoured by various cosmological constraints, as we will see later, we neglect the constraints from stellar cooling in our study.
Long-lived ALPs that survive after recombination may still be observable today through their decay products. The photons injected by ALP decays are the most detectable candidates, as they can be probed in observations of the extra-galactic background light (EBL) or via X-rays [41]. For ALP masses below the keV scale, the constraints on the ALP-photon coupling from X-ray and EBL surveys require that $g_{a\gamma}\lesssim 10^{-15}$ GeV-1 [41, 44]. In our model, however, the ALP-photon coupling is derived at two-loop order, and therefore extremely small. As above, only for very large mixing $U_{eN}^{2}\sim 10^{-3}$ and large HNL masses $m_{N}\sim 1$ TeV, has this constraint any effect. Consequently, we also ignore any astrophysical constraints from late-time photonic decays of ALPs.
@@ 4.5 Impact on direct HNL searches
Future HNL searches, especially those based on long-lived signatures, will probe small active-sterile mixing strengths approaching the seesaw expectation. Prominent examples of proposed and planned searches are PIONEER [82], NA62 [83, 84], DUNE [85], SHiP [86], FCC-ee [87]. A dedicated analysis is needed to present a study of the sensitivity in our scenario depending on the specific search strategy. Such a detailed analysis is beyond the scope of this paper and will be considered as future work. In this section, we will assess how the sensitivity is modified in presence of new HNL-ALP coupling in the context of DUNE. We qualitatively describe the approach to calculate the expected number of signal events in the DUNE near detector (ND) based on the analysis [88]. The ND is located at a distance of $L=574$ m from the HNL production point, with a transverse cross-section of $A=12~{}\rm{m}^{2}$ and a depth of $\Delta L=5$ m along the beam axis. Following [88], we write the expected number of signal events as
$N_{\rm{sig}}=N_{P}\times\rm{Br}(P\rightarrow N)\times\rm{Br}(N\rightarrow\rm{charged})\times\epsilon_{\rm{geo}},$ (4.3)
@@
$\epsilon_{\rm{geo}}=e^{-\frac{m_{N}\Gamma_{N}}{p_{N_{z}}}L}\left(1-e^{-\frac{m_{N}\Gamma_{N}}{p_{N_{z}}}\Delta L}\right).$ (4.4)
Here, $p_{N_{z}}$ is the momentum of the HNL along the beam axis in the lab frame, where we have considered $p_{N_{z}}=7.5$ GeV, following the simulation of meson production from a $pp$ collision at $\sqrt{s}=15$ GeV [88]. As we have not explicitly simulated any events and analysed with respect to detector level cuts, for simplicity we will consider, in our analysis, that all produced HNL events will be accepted at detector level.
@@
$\frac{N^{\prime}_{\rm{sig}}}{N_{\rm{sig}}}=\frac{\rm{Br^{\prime}}(N\rightarrow\rm{charged})}{\rm{Br}(N\rightarrow\rm{charged})}\frac{\epsilon^{\prime}_{\rm{geo}}}{\epsilon_{\rm{geo}}}$ (4.5)
@@
$\frac{\epsilon^{\prime}_{\rm{geo}}}{\epsilon_{\rm{geo}}}=\exp\left[-\frac{m_{N}}{p_{N_{z}}}\Gamma(N\to a\nu)L\right]\frac{\Gamma^{\prime}_{N}}{\Gamma_{N}},$ (4.6)
for a shallow detector depth $\Delta L$, i.e., for $(m_{N}/p_{N_{z}})\Gamma^{\prime}_{N}\Delta L\ll 1$. The ratio of decay rates in this expression is equal to $\Gamma^{\prime}_{N}/\Gamma_{N}=\text{Br}(N\to\text{charged})/\rm{Br^{\prime}}(N\to\text{charged})$, thus cancelling the corresponding ratio in Eq. (4.5) in this limit. We thus have
$\frac{N^{\prime}_{\rm{sig}}}{N_{\rm{sig}}}=\exp\left[-\frac{m_{N}}{p_{N_{z}}}\Gamma(N\to a\nu)L\right],$ (4.7)
@@
As another example for a direct search experiment we consider the existing NA62 experiment which uses a different search strategy as compared to DUNE. The NA62 experiment [83] used a secondary 75 GeV hadron beam containing a fraction of kaons, and has been able to probe the decays $K^{+}\to\ell^{+}N$. For small active-sterile mixing the decay length of HNL is much larger than the 75 m detector size and the process is characterised by a single detected track, that of the charged lepton – a positive signal is a peak in its missing mass distribution. As this experiment is currently insensitive to the decay of the HNLs, the presence of the $N\to a\nu$ decay channel will not affect the search as long as the decay length remains large.
@@ 5 Results and discussion
The aim of this work is to find valid scenarios where the traditional BBN bound on HNLs [47] is relaxed due to the primary decays to ALPs. For this purpose we have performed some small scale parameter scans around the benchmark scenarios in Table 1 to illustrate this effect [^3] We performed a grid scan on two parameters of the model, $m_{N}$ and $|U_{eN}|^{2}$. This simplification is enough for our purposes as we only intend to highlight the crucial features of the model, but do not attempt to map the available parameter space nor perform any statistical interpretation of the results. For more details on the rigorous treatment of inference and statistics in physics see [89].. Figure 6 shows the results of these parameter scans, varying the values of $m_{N}=[10^{-3},10^{2}]$ GeV and $|U_{eN}|^{2}=[10^{-13},10^{-6}]$. The top row corresponds to scans around scenarios 1 and 2, with only $m_{a}$ and $f_{a}$ fixed, and, where the blue diamond markers are the fixed values of $m_{N}$ and $|U_{eN}|^{2}$ from Table 1. The bottom row shows scans of scenarios 3 and 4, which have a lighter ALP mass of $m_{a}=10$ eV.
@@
The strongest cosmological constraint, shown as the shaded brown region in the panels, arises from the increase on the neutrino temperature due to ALP decays, which shows as a modification of the $N_{\rm eff}$ at the time of recombination. $N_{\rm eff}$ constrains the larger HNL masses and mixing, corresponding to shorter ALP lifetimes. Consequently, ALPs decaying before the onset of BBN, or shortly after, would cause a sufficient increase of the neutrino temperature and be excluded by the Planck measurement of the $N_{\rm eff}$ value. Smaller HNL-ALP couplings cause longer ALP lifetimes, and thus the right-handed panels with larger $f_{a}$ have weaker $N_{\rm eff}$ constraints. However, for the bottom panels with lower $m_{a}$, even though the lifetime is also longer and one would expect the limit to be weaken, the opposite effect occurs and the $N_{\rm eff}$ limit is significantly stronger. This happens because lighter ALPs are ultra-relativistic for a longer period between BBN and recombination, and thus have a stronger effect on the expansion rate, and thus on the neutrino temperature and value of $N_{\rm eff}$. Other cosmological constraints such as, energy injection before BBN or the relic abundance of ALPs are subleading and thus not shown. The region above the purple contour is, in principle, excluded due to the astrophysical constraints from supernova SN1987A, arising from the modified neutrino flux due to the production of HNLs in the core of the supernova. Nevertheless, the uncertainties in the calculation of the SN neutrino flux are high and depend on the choice of SN core model [90], hence we show in purple a conservative expectation of the limit, but refrain from excluding models only on the basis of this SN constraint. Supernova constraints due to ALP production, which is largely independent of $m_{a}$ in the ranges of interest, only apply for large values of $|U_{eN}|^{2}$ and $m_{N}$, which however resides well outside of our considered region.
@@
In Fig. 7 we present this newly available parameter region which may be the target of future dedicated searches for HNLs. In the left panel, we show the corresponding parameter space for scenario 1 ($f_{a}=1$ TeV, $m_{a}=1$ keV) as described in previous section, while the right panel corresponds to the scenario 2 ($f_{a}=10^{2.5}$ TeV, $m_{a}=1$ keV). The gray shaded regions in both panels are disfavored from various cosmological and astrophysical constraints discussed in Sec. 4, and shown coloured in Figure 6. The gray dashed curve denotes the BBN bound from the HNL only decaying to SM decay channels i.e., in absence of the ALP. The corresponding seesaw regime is shown in blue considering $1~{}\rm{meV}~{}\lesssim m_{\nu}\lesssim 100~{}\rm{meV}$. The red dashed contour corresponds to the future sensitivity of DUNE HNL searches in standard HNL models (without ALPs), while the solid red contours in both panels denote the resultant sensitivity contours for observation of 6 events in presence of the new $N\rightarrow a\nu$ decay channel in respective benchmark scenarios with two different $f_{a}$ values. Additionally, as described in Tab. 1, in blue diamonds two benchmark scenarios with fixed HNL mass and mixing $m_{N}=10^{-1}$ GeV, $|U_{eN}|^{2}=10^{-10}$ and $m_{N}=10^{-0.4}$ GeV, $|U_{eN}|^{2}=10^{-9.2}$, have been shown. The brown dashed contour denotes the present exclusion contour from NA62 experiment. Furthermore, considering the discussion of subsection 4.5, we have also highlighted in brown the region where the approximation holds (with $L_{\rm{NA}62}$ = 75 m and $p_{N_{z}}\sim$ 30 GeV as the $K^{+}$ beams have energy $\sim 75$ GeV [83]), the present sensitivity contour remains unaffected by the new axion decay channel of HNL in this region. From Figure 7 it is evident that the available regions constrained by several astrophysical and cosmological constraints can be well probed in the future by DUNE and NA62 in presence of this new decay channel of HNLs.